OpenStudy (riley):

Calculate dy/dx. Simplify your answer. y = x^5√x

OpenStudy (riley):

I know I use the power rule, after multiplying it out. But the square root is confusing me.

OpenStudy (anonymous):

\[x ^{5}\sqrt{x}=x^{5+1/2}=x^{5.5}\]

OpenStudy (anonymous):

Hence \[dx/dy= 5.5x^{4.5}=5.5x^4\sqrt{x}\]

OpenStudy (anonymous):

I assume you know \[\sqrt{x}=x^{1/2}\]

OpenStudy (riley):

so you just multiplied x^5 and x^1/2. And then used the power rule, right? O_O

OpenStudy (riley):

Yes I knew that. x3

OpenStudy (anonymous):

Exactly

OpenStudy (riley):

Ohhhh okay, I see I see. That square root was tripping me up so badly. thank you very much for your help.

OpenStudy (anonymous):

you would use the product rule because it's technically (x^5)(x^1/2)

OpenStudy (anonymous):

either way.